Do heavier people burn more energy? A researcher collected data on the lean body mass (LBM) and resting metabolic rate (MR) for 10 individuals who were subjects of the study of dieting. LBM given in kilograms, is person’s weight leaving out all fat. Metabolic rate (MR) measured in calories burned per 24 hours. The researcher believes that LBM (independent) is an important influence on MR (dependent). Scores for variables body mass and metabolic rate are given in the following table (the scores have been rounded). The MR scores are in 10s scale. Write the rest of column headers and finish the regression table in the process as you are answering the following questions.

A. Obtain the slope (b) and the y-intercept (a) of the least square regression.

B. Write the regression equation for the body mass (X) and metabolic rate (Y) and identify b and a in the equation.

C. Predict the value of metabolic rate if body mass would be 112

D. Use correlation formula to calculate R , the relationship between body mass and metabolic rate

E. What does R2 tell us about the relationship between the two variables?

Body Mass (X) Metabolic Rate (Y)   \overline{X} - X\overline{Y} - Y

40 100

56 140

49 141

40 142

51 151

33 95

43 111

52 135

42 121

34 104
Answer #1

Tiven that depende ut [ İndetendent. siable tboo yovo too l40 lu l 14 2 Luo S6 49 11o s1 7840 3/3 6 61 24/ S68。 )/600 77026 B3 95 4773 o uq 2764 3536 16 S670 1OO 13厂 1 21 УЗ S 2 4 2 0440 19900 2 y = 1240-124paae Thus the esesin sp(b 19100 1ox u 2-0666 ano = 124-2.0667x44 33 . 066구 Soluhem (BPas の7 . x= 112. el イニ33.067 + 2.067 x112 ープ) | Y-264.533. e D Soluho 240-4 ,toV . :.J讶: 76 6 12. o16 92 144 25 2201 49 to 12 72 l &G 3/9 329 49 72 121 ) 1 9 93 13 13 52 & 135 11 y2-2121-3 ㄋㄐ -10 2 - 20 190 20。 4 6 314 /240 Solu bon ekindly give your opinion